Shift Summary, June 17, 1999 Evening

Peter Shanahan (shanahan@fnal.gov)
Fri, 18 Jun 1999 00:18:09 -0500

Shift Summary for Eve shift, 17/Jun/99

On Shift: Patty McBride, Peter Shanahan

Hours of beam delivered at all intensities: 7.75
Hours of beam delivered at nominal intensity: 7.5
Hours of E832 data (special runs not included): 7.00
Hours of special runs (laser scans, muon runs...): 0.75

Accounting of lost time: A few minutes due to LINAC problems

========================================================================
Runs ended this shift:
Run 13638 - 15.1K offline VAC00 - BEAM832
Run 13639 - 11.2K offline VAC00 - BEAM832
Run 13640 - ~.5M mu - MUON DC1,2 window HV=2200V
Run 13641 - ~.5M mu - MUON DC1,2 window HV=2400V (nominal)
Run 13642 - ~.3M mu - MUON NM4AN @ 100 MeV/c for mu spectrum

Tapes brought to FCC this shift: None
(including tapes for online split)

========================================================================
Accesses during shift: None

========================================================================
Overview of shift:
Normal data taking. Increase beam request to 5.E12 prior to 13639.
This was met according to our SEM, but total TeV extraction reads
~4.2E12.

Muon runs included DC1/2 HV studies, and mu energy spectrum data taking.

========================================================================
Description of Shift:

run spill time comment
------------------------------------------------------------------------
13638 16:04 NM1U reads back ~200 A low for 1 second every
several spills. Rick F and MCR investigate:
probably just a readback problem.

13638 19:00 Dave N installed 3 new Ar-E bottles on west side

13639 1 19:12 Ed changed B06 (3pi0) prescale from 1 to 2.
13639 1 19:12 Request changed to 5E12. MCR complies almost
immediately
13639 ~40-57 19:12- Beam was wandering by +/- 1mm in both directions at
19:29 target. MCR was aware of the problem, and fixed it.

13639 20:37 Called MCR to complain about intensity ~5.35E12.
They say that real intensity is much lower, as
indicated by total TeV FT ~ 4.3E12

13640 1 23:09 Set up for muon running. DC12 window HV modified.
13641 1 23:24 Muon run, DC12 window wire HV nominal.
13642 1 23:38 Muon run, but with NM4AN @ 100 GeV/c

13643 1 23:55 Return to nominal conditions, with NM4AN=-1600 A.
13643 2 23:56 Realize that we forgot to pull out beam stop. Fixed.

========================================================================
Follow up Notes on Previous Shift

No Work done on NM2EU1.

========================================================================
Items to Follow up on Next Shift

Tomorrow morning after 9am there will be radiation survey in KTeV
enclosure for 1-2 hours. (We maybe able to run some muons with sweeper
magnet set high)